COMPLETE THE ORDERED PAIRS:
2X+3Y=7
(0, ) (2, ) ( ,3)

Answer by oscargut(2103)

2X+3Y=7
2X+3Y=7
(0, ) (2, ) ( ,3)
if x=0 then 3y=7 then y=7/3
if x=2 then 4+3y=7 then 3y=3 then y=1
if y=3 then 2x+9=7 then 2x=-2 then x=-1
so the pairs are:
(0,7/3) (2,1) (-1,3)